# Constraints on the Variation of the QCD Interaction Scale $\Lambda_{\text{QCD}}$

### v3

@@ IV Constraint from the Oklo natural nuclear reactor
where $E_{k}$ is the kinetic energy of the captured neutron inside the nucleus and $C_{i}$ are dimensionless parameters. The contribution of the Coulomb interaction is proportional to $e^{2}/r_{0}\propto{\alpha}\Lambda_{\text{QCD}}$, see Eq. (16). Substituting the expressions for $E_{k}$ and $U$ into the resonance energy gives
@@
The shift of the resonance due to the variation of $\alpha$ was calculated in Refs. [72, 73] to be
@@
was calculated in Ref. [26]. Using these results, we obtain
@@
Now we should specify how the shift of $E_{r}$ has been measured [72, 73]. It was expressed through the measured ratio of abundances of 149Sm and 150Sm near the Oklo natural nuclear reactor and compared to this ratio in other areas where there was no neutron flux transforming 149Sm to 150Sm. The neutron flux in Oklo was produced by the nuclear fission of 235U, and the integrated neutron flux was found by comparing the difference between the abundance of 235U in Oklo and its natural abundance in other places.
The observed shift in the resonance energy $\delta E_{r}$ between the present and time of the Oklo natural reactor 1.8 billion years ago was constrained to be less than 0.02 eV [75, 76, 74], which provides the constraint
@@
This result can be interpreted in the context of dark-energy-type models [20, 21, 22, 23, 1], where variations occur on cosmological scales. Another class of models involves dark halos, in which the dark matter density around massive bodies is strongly enhanced compared to the average dark matter density and evolves on timescales of billions of years (see, e.g., [77, 78, 79]). Dynamical mechanisms of this type predict present-day fractional variations at the level of $10^{-13}$$10^{-16}$ [66], far below the $10^{-9}$ sensitivity accessible with Oklo. Therefore, Oklo may only probe an exceptionally large increase of the dark matter density. Large variations in the dark matter density could also accur due to passing dark matter clouds or Bose stars. In this context, Oklo constraints on $\Lambda_{\text{QCD}}$ may rule out the possibility that such extreme local enhancements of the dark matter density occurred in the past or are occurring today.
@@
V Constraints from Quasar Absorption Spectra
Quasar absorption spectra provide an extremely sensitive astrophysical probe of the variation of fundamental constants over cosmological timescales - see e.g [20, 21, 22, 23, 1, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96]. Intervening clouds along the line of sight of the quasar give access to the spectra of the atoms and ions present in the cloud. Calculations of the dependence of corresponding atomic transition frequencies on the fundamental constants have been performed e.g. in Refs. [97, 98, 14, 15, 99, 100, 101] and references therein. Through the analysis of atomic and molecular transitions in such absorption systems, measurements of the variation of the fine structure constant $\alpha$, the proton to electron mass ratio $\mu$ and composite parameters $F\equiv g_{p}(\alpha^{2}\mu)^{1.57}$, $x\equiv\alpha^{2}g_{p}/\mu$, $F^{\prime}\equiv\alpha^{2}\mu$ and $G\equiv g_{p}(\alpha\mu)^{1.85}$ have been performed, where $g_{p}$ is the proton $g$-factor; see e.g. Refs. [82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96].
@@ V Constraints from Quasar Absorption Spectra
$\mu$ $(2.78\pm 0.88)$ 2.59 [82] 1.12 $(2.78\pm 0.88)$ $2.49\pm 0.79$
$(2.06\pm 0.79)$ 3.02 [82] 1.13 $(2.06\pm 0.79)$ $1.78\pm 0.68$
$(1.01\pm 0.62)$ 2.59 [83] 1.12 $(1.01\pm 0.62)$ $0.90\pm 0.55$
$(0.82\pm 0.74)$ 2.8 [83] 1.14 $(0.82\pm 0.74)$ $0.72\pm 0.65$
$(0.26\pm 0.30)$ 3.02 [83] 1.16 $(0.26\pm 0.30)$ $0.22\pm 0.26$
$(0.70\pm 0.80)$ 3.02 [84] 1.16 $(0.70\pm 0.80)$ $0.61\pm 0.69$
$(0.56\pm 0.55_{\text{stat}}\pm 0.27_{\text{syst}})$ 2.06 [85] 1.05 $(0.56\pm 0.61)$ $0.53\pm 0.58$
<0.18<0.18 0.685 [86] 0.629 <0.18<0.18 <0.29<0.29
<0.38<0.38 0.685 [87] 0.629 <0.38<0.38 <0.60<0.60
<0.14<0.14 0.89 [88] 0.638 <0.14<0.14 <0.22<0.22
<9<9 2.42 [89] 1.10 <9<9 <8.2<8.2
$F$ $(-0.44\pm 0.36\pm 1.0_{\text{syst}})$ 0.765 [90] 0.673 $(-0.28\pm 0.68)$ $0.42\pm 1.0$
$(0.51\pm 1.26)$ 0.247 [91] 0.293 $(0.33\pm 0.80)$ $1.11\pm 2.74$
$x$ $(-0.63\pm 0.99)$ 1.29 [92] 0.883 $(0.63\pm 0.99)$ $0.71\pm 1.12$
$(-0.17\pm 0.17)$ 3.17 [93] 1.17 $(0.17\pm 0.17)$ $0.15\pm 0.15$
$F^{\prime}$ $(1.0\pm 10)$ 4.69 [94] 1.24 $(1.00\pm 10.00)$ $0.80\pm 8.0$
$(14\pm 15)$ 6.42 [94] 1.29 $(14\pm 15)$ $11\pm 12$
$G$ <1.1<1.1 0.247 [95] 0.294 <0.60<0.60 <2.03<2.03
<1.16<1.16 0.0018 [95] 0.0025 <0.63<0.63 <250<250
$(-1.18\pm 0.46)$ 0.247 [96] 0.294 $(-0.64\pm 0.25)$ $2.17\pm 0.85$
Table 3: Constraints on the variation of $\Lambda_{\text{QCD}}$ from Quasar Absorption Spectra. The constants (as presented in Ref. [1]) are: $\mu\equiv m_{p}/m_{e}$, $F\equiv g_{p}(\alpha^{2}\mu)^{1.57}$, $x\equiv\alpha^{2}g_{p}/\mu$, $F^{\prime}\equiv\alpha^{2}\mu$ and $G\equiv g_{p}(\alpha\mu)^{1.85}$.
The proton mass $m_{p}$ and magnetic $g$-factor $g_{p}$ both depend on $\Lambda_{\text{QCD}}$. Within 10% accuracy, $\frac{\delta\mu}{\mu}\approx\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}$ - see Eq. (20). A detailed calculation of the quark mass dependence of $g_{p}$ in Ref. [14] together with Eqs. (5,6) leads to the estimate $\frac{\delta g_{p}}{g_{p}}\approx 0.10\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}$. This allows us to translate existing astrophysical bounds, as compiled in Ref. [1], into direct constraints on $\Lambda_{\text{QCD}}$, see Table 3. Under the assumption of a linear variation of $\Lambda_{\text{QCD}}$, these constraints are remarkably close to those obtained from atomic clock measurements. We note, however, that this assumption has no justification over such an extensive period of time.
VI Summary
@@
Within the minimal scenario in which this gluonic channel is the sole $\phi$ interaction, we show that the atomic Yb/Cs clock comparison [4] places the constraint $\dot{\Lambda}_{\text{QCD}}/\Lambda_{\text{QCD}}=(3.2\pm 3.5)\times 10^{-17}\ \text{yr}^{-1}$. Further independent constraints are obtained from isotopic ratios measured in the Oklo natural nuclear reactor, which limit the variation to $|\delta\Lambda_{\text{QCD}}/\Lambda_{\text{QCD}}|<2\times 10^{-9}$ over 1.8 billion years. Assuming a linear drift, this translates to a bound $|\dot{\Lambda}_{\text{QCD}}/\Lambda_{\text{QCD}}|<1\times 10^{-18}\text{yr}^{-1}$. Additionally, the sensitivity of the 229Th nuclear clock to the variation of $\Lambda_{\text{QCD}}$ is estimated to be four orders of magnitude higher than that of current atomic clocks.

### v4

@@ V Constraint from the Oklo natural nuclear reactor
where $E_{k}$ is the kinetic energy of the captured neutron inside the nucleus and $C_{i}$ are dimensionless parameters. The contribution of the Coulomb interaction is proportional to $e^{2}/r_{0}\propto{\alpha}\Lambda_{\text{QCD}}$. Substituting the expressions for $E_{k}$ and $U$ into the resonance energy gives
@@
The shift of the resonance due to the variation of $\alpha$ was calculated in Refs. [Shlyakhter1976, DAMOUR1996] to be
@@
was calculated in Ref. [Wiringa2009]. Using these results, we obtain
@@
Now we should specify how the shift of $E_{r}$ has been measured [Shlyakhter1976, DAMOUR1996]. It was expressed through the measured ratio of abundances of 149Sm and 150Sm near the Oklo natural nuclear reactor and compared to this ratio in other areas where there was no neutron flux transforming 149Sm to 150Sm. The neutron flux in Oklo was produced by the nuclear fission of 235U, and the integrated neutron flux was found by comparing the difference between the abundance of 235U in Oklo and its natural abundance in other places.
The observed shift in the resonance energy $\delta E_{r}$ between the present and time of the Oklo natural reactor 1.8 billion years ago was constrained to be less than 0.02 eV [Gould2006, Petrov2006, FUJII2000], which provides the constraint
@@
This result can be interpreted in the context of dark-energy-type models [BarrowMagueijo1998, Barrow2005, BarrowSandvikMagueijo2002, Wetterich2003, Uzan], where variations occur on cosmological scales. Another class of models involves dark halos, in which the dark matter density around massive bodies is strongly enhanced compared to the average dark matter density and evolves on timescales of billions of years (see, e.g., [BanerjeeFlambaum2020, Banerjee2020, Budker2023]). Dynamical mechanisms of this type predict present-day fractional variations at the level of $10^{-13}$$10^{-16}$ [Fuchs2024PRX], far below the $10^{-9}$ sensitivity accessible with Oklo. Therefore, Oklo may only probe an exceptionally large increase of the dark matter density. Large variations in the dark matter density could also accur due to passing dark matter clouds or Bose stars. In this context, Oklo constraints on $\Lambda_{\text{QCD}}$ may rule out the possibility that such extreme local enhancements of the dark matter density occurred in the past or are occurring today.
@@
VI Constraints from Quasar Absorption Spectra
Quasar absorption spectra provide an extremely sensitive astrophysical probe of the variation of fundamental constants over cosmological timescales - see e.g [BarrowMagueijo1998, Barrow2005, BarrowSandvikMagueijo2002, Wetterich2003, Uzan, Murphy2001, MurphyWebbFlambaum2003, Reinhold2006, King2008, Thompson2009, Malec2009, Murphy2008, Menten2008, Henkel2009, Petitjean2009, Kanekar2005, Darling2004, Tzanavaris2007, Srianand2010, Levshakov2008, Chengalur2003, Kanekar2010]. Intervening clouds along the line of sight of the quasar give access to the spectra of the atoms and ions present in the cloud. Calculations of the dependence of corresponding atomic transition frequencies on the fundamental constants have been performed e.g. in Refs. [DZUBAPRL1999, DZUBAPRA1999, Thomas, Tedesco, CanJPhQ, Ammonia, KozlovMolecules] and references therein. Through the analysis of atomic and molecular transitions in such absorption systems, measurements of the variation of the fine structure constant $\alpha$, the proton to electron mass ratio $\mu$ and composite parameters $F\equiv g_{p}(\alpha^{2}\mu)^{1.57}$, $x\equiv\alpha^{2}g_{p}/\mu$, $F^{\prime}\equiv\alpha^{2}\mu$ and $G\equiv g_{p}(\alpha\mu)^{1.85}$ have been performed, where $g_{p}$ is the proton $g$-factor; see e.g. Refs. [Reinhold2006, King2008, Thompson2009, Malec2009, Murphy2008, Menten2008, Henkel2009, Petitjean2009, Kanekar2005, Darling2004, Tzanavaris2007, Srianand2010, Levshakov2008, Chengalur2003, Kanekar2010].
The proton mass $m_{p}$ and magnetic $g$-factor $g_{p}$ both depend on $\Lambda_{\text{QCD}}$. Within 10% accuracy, $\frac{\delta\mu}{\mu}\approx\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}$ - see Eq. (12). A detailed calculation of the quark mass dependence of $g_{p}$ in Ref. [Thomas] together with Eqs. (5,6) leads to the estimate $\frac{\delta g_{p}}{g_{p}}\approx 0.10\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}$. This allows us to translate existing astrophysical bounds, as compiled in Ref. [Uzan], into direct constraints on $\Lambda_{\text{QCD}}$, see Table 3. Under the assumption of a linear variation of $\Lambda_{\text{QCD}}$, these constraints are remarkably close to those obtained from atomic clock measurements. We note, however, that this assumption has no justification over such an extensive period of time.
VII Summary
@@
$\mu$ $(2.78\pm 0.88)$ 2.59 [Reinhold2006] 1.12 $(2.78\pm 0.88)$ $2.49\pm 0.79$
$(2.06\pm 0.79)$ 3.02 [Reinhold2006] 1.13 $(2.06\pm 0.79)$ $1.78\pm 0.68$
$(1.01\pm 0.62)$ 2.59 [King2008] 1.12 $(1.01\pm 0.62)$ $0.90\pm 0.55$
$(0.82\pm 0.74)$ 2.8 [King2008] 1.14 $(0.82\pm 0.74)$ $0.72\pm 0.65$
$(0.26\pm 0.30)$ 3.02 [King2008] 1.16 $(0.26\pm 0.30)$ $0.22\pm 0.26$
$(0.70\pm 0.80)$ 3.02 [Thompson2009] 1.16 $(0.70\pm 0.80)$ $0.61\pm 0.69$
$(0.56\pm 0.55_{\text{stat}}\pm 0.27_{\text{syst}})$ 2.06 [Malec2009] 1.05 $(0.56\pm 0.61)$ $0.53\pm 0.58$
<0.18<0.18 0.685 [Murphy2008] 0.629 <0.18<0.18 <0.29<0.29
<0.38<0.38 0.685 [Menten2008] 0.629 <0.38<0.38 <0.60<0.60
<0.14<0.14 0.89 [Henkel2009] 0.638 <0.14<0.14 <0.22<0.22
<9<9 2.42 [Petitjean2009] 1.10 <9<9 <8.2<8.2
$F$ $(-0.44\pm 0.36\pm 1.0_{\text{syst}})$ 0.765 [Kanekar2005] 0.673 $(-0.28\pm 0.68)$ $0.42\pm 1.0$
$(0.51\pm 1.26)$ 0.247 [Darling2004] 0.293 $(0.33\pm 0.80)$ $1.11\pm 2.74$
$x$ $(-0.63\pm 0.99)$ 1.29 [Tzanavaris2007] 0.883 $(0.63\pm 0.99)$ $0.71\pm 1.12$
$(-0.17\pm 0.17)$ 3.17 [Srianand2010] 1.17 $(0.17\pm 0.17)$ $0.15\pm 0.15$
$F^{\prime}$ $(1.0\pm 10)$ 4.69 [Levshakov2008] 1.24 $(1.00\pm 10.00)$ $0.80\pm 8.0$
$(14\pm 15)$ 6.42 [Levshakov2008] 1.29 $(14\pm 15)$ $11\pm 12$
$G$ <1.1<1.1 0.247 [Chengalur2003] 0.294 <0.60<0.60 <2.03<2.03
<1.16<1.16 0.0018 [Chengalur2003] 0.0025 <0.63<0.63 <250<250
$(-1.18\pm 0.46)$ 0.247 [Kanekar2010] 0.294 $(-0.64\pm 0.25)$ $2.17\pm 0.85$
Table 3: Constraints on the variation of $\Lambda_{\text{QCD}}$ from Quasar Absorption Spectra. The constants (as presented in Ref. [Uzan]) are: $\mu\equiv m_{p}/m_{e}$, $F\equiv g_{p}(\alpha^{2}\mu)^{1.57}$, $x\equiv\alpha^{2}g_{p}/\mu$, $F^{\prime}\equiv\alpha^{2}\mu$ and $G\equiv g_{p}(\alpha\mu)^{1.85}$.
@@
Within the minimal scenario in which this gluonic channel is the sole $\phi$ interaction, we show that the atomic Yb/Cs clock comparison [YbCs] places the constraint $\dot{\Lambda}_{\text{QCD}}/\Lambda_{\text{QCD}}=(3.2\pm 3.5)\times 10^{-17}\ \text{yr}^{-1}$. Further independent constraints are obtained from isotopic ratios measured in the Oklo natural nuclear reactor, which limit the variation to $|\delta\Lambda_{\text{QCD}}/\Lambda_{\text{QCD}}|<2\times 10^{-9}$ over 1.8 billion years. Assuming a linear drift, this translates to a bound $|\dot{\Lambda}_{\text{QCD}}/\Lambda_{\text{QCD}}|<1\times 10^{-18}\text{yr}^{-1}$. Additionally, the sensitivity of the 229Th nuclear clock to the variation of $\Lambda_{\text{QCD}}$ is estimated to be four orders of magnitude higher than that of current atomic clocks.